Fermi National Laboratory
 

Accelerator Update for October 25, 2006 - October 27, 2006

 

Wednesday October 25

The day shift began with Operations monitoring the Tevatron (TeV) store 5026 (60.24E30) and the Recycler stash (82.5E10), with the Antiproton Source (Pbar) stacking (44.2E10), with Linac, Booster, and the Main Injector (MI) operating normally, with MiniBooNE and NuMI taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

At 9:02 AM, a Booster power supply (WAPS) tripped off with no indication of the problem and wouldn't reset. The Booster 400MeV line valves were closed, and RF station (BRF15) was down. Operators also discovered that the Tevatron Electron Lens (TEL1) had tripped, NuMI went down, and a MI RF station (MIRF3) had tripped.

At 9:11 AM, a high voltage expert reported a ComEd glitch had occurred at 8:59 AM. This was a voltage sag of short duration on a 345 KV line that feeds the Master Substation.

A MI expert began a MI-8 study at 9:53 AM.

The Recycler completed a stash of antiprotons at 12:14 PM.

The Recycler completed a stash around 2:54 PM.

At 4:49 PM, an expert began work on a Debuncher power supply (D:IB).

The Recycler completed a stash around 5:15 PM.

At 5:15 PM, Operations reported that the current from the H- Source had been erratic.

The work on D:IB ended and stacking resumed at 5:29 PM.

At 6:14 PM, Operations switched from the H- Source to the I- Source. An expert came in and tuned the new source.

Operations terminated store 5026 at 9:02 PM.

Operations prepared CDF for access at 9:12 PM.

Operations reported at 10:22 PM that the I- Source output was at 39 mA, which was higher than normal.

Operations put the TeV through a dry squeeze at 10:23 PM

Operations put the TeV into shot setup at 11:11 PM.

At 11:19 PM, a Debuncher power supply (D:QD) tripped off on a ground fault indication. It reset okay, but this supply had the same problem a few days ago; operators watched it closely.

Thursday October 26

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (251.1E10), with Pbar stacking (32.3E10), and with NuMI and MiniBooNE taking beam.

At 12:27 AM, Operations reported that D:QD had already tripped off twice. It went on to trip off four more times over the next seven hours. It reset each time.

At 2:52 AM, Operations reported that they had lost one transfer of antiprotons from the Recycler to the TeV due to a Recycler RF problem.

Operations established store 5028 at 3:36 AM with an initial luminosity of 156.4E30.

At 9:16 AM, Operations prepared Pbar for an access so experts could search for a ground fault at D:QD. The expert reported at 10:17 AM that he hadn't found a fault.

Pbar resumed stacking at 11:35 AM.

The Neutron Therapy Facility (NTF) reported at 2:09 PM that they had finished treating patients for the day.

The Recycler completed a stash around 3:18 PM.

At 4:33 PM, store 5028 aborted and the TeV quenched at sectors B1 and D1. This was again due to the loss of the F sector audio beam permit. Safety system technicians investigated.

Operations prepared CDF for an access at 5:26 PM.

Operations reported at 6:04 PM that the safety system techs had found a faulty actuator for the F Sector key #31 switch. They had the actuator replaced an hour later.

NuMI, MiniBooNE, and Pbar all resumed taking beam by 7:18 PM.

Operations put the TeV through a dry squeeze at 8:19 PM.

At 8:25 PM, a Recycler power supply (R:C301) tripped off. After consulting an expert, Operations reset the supply.

Operations put the TeV into shot setup at 8:43 PM.

At 9:40 PM, Operations reported stochastic cooling PLC errors. Operators contacted experts.

Operations established store 5030 at 10:31 PM with an initial luminosity of 132.53E30.

At 10:56 PM, Operations reported that the TeV D17 separator had sparked. This caused a loss of luminosity.

At 11:20 PM, operators reported that they had followed the Recycler experts' suggestion of rebooting the PLCs in MI-11 and MI-21. Multiple reboots did not work and neither did a power cycle. Operators tried to contact PLC experts. This problem halted stashing.

Friday October 27

The midnight shift began with Operations monitoring store 5030 (98.7E30) and the stash (2.8E10), with the Recycler unable to stash due to a PLC problem, with Pbar stacking (101.4E10), and with NuMI and MiniBooNE taking beam.

At 12:46 AM, Operations reported that they were still trying to find an expert to help them with the Recycler PLC problem. Two people who knew something of the system, but weren't experts agreed to come in and help troubleshoot. At 3:49 AM, they reported a possible problem with the VME.

Operations reported at 5:29 AM that a Recycler expert had come in and found a spare VME card. The expert said he replaced the card, rebooted various systems, and the system control returned to normal, but he wasn't sure that he had fixed the problem.

Operations reported at 5:54 AM that they were just starting their third Recycler stash.

The Plans for Friday and the Weekend

The plan for today is stack and stash, and allow some studies. The Run Coordinator has not decided when the next store would go in, but possibly late on the evening shift or early on the midnight shift. The weekend will consist of stacking, stashing, and storing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/27/2006    email Fermilab